# Omnimodal Understanding for Multimodal LLMs

### v2

@@ 4.3 Roadmap Towards Real-world Understanding
[…]
@@
[…]
@@
Methods Speech Event Music Overall Video + Subtitle + Audio Video + Subtitle + Audio Video + Subtitle + Audio Video + Subtitle + Audio Unified-IO-2 L [43] 26.8 13.9-12.9 23.1-3.1 26.9 13.5-13.4 23.2-3.7 26.3 15.0-11.3 24.5-1.8 26.6 14.8-11.8 23.3-3.3 Unified-IO-2 XL [43] 25.0 13.0-12.0 24.6-0.4 24.8 12.3-12.5 25.0+0.2 26.7 15.9-10.8 24.7-2.0 25.3 14.1-11.2 24.7-0.6 Unified-IO-2 XXL [43] 27.0 15.6-11.4 25.7-1.3 26.2 14.2-12.0 25.1-1.1 28.4 19.1-9.3 25.9-2.5 27.2 17.2-10.0 25.9-1.3 OneLLM [22] 12.5 19.6+7.1 23.2+10.7 12.4 19.3+6.9 22.3+9.9 12.4 19.0+6.6 22.5+10.1 12.6 19.6+7.0 22.8+10.2 VideoLLaMA2 [7] 17.1 25.5+8.4 25.2+8.1 16.1 24.9+8.8 24.5+8.4 17.7 27.0+9.3 25.6+7.9 17.4 26.1+8.7 25.4+8.0 VITA-1.5 [18] 37.6 39.1+1.5 36.2-1.4 36.4 38.2+1.8 35.3-1.1 38.7 40.0+1.3 37.1-1.6 37.7 39.3+1.6 36.5-1.2 Qwen2.5-Omni [71] 38.7 38.7+0.0 44.8+6.1 37.6 37.7+0.1 43.8+6.2 40.7 40.3-0.4 46.1+5.4 39.2 39.2+0.0 45.2+6.0 Gemini 1.5 Pro [60] 34.3 39.6+5.3 49.2+14.9 33.0 38.9+5.9 48.7+15.7 35.4 39.2+3.8 49.8+14.4 34.4 39.3+4.9 48.0+13.6
@@
Methods Speech Event Music Overall Video + Subtitle Video + Subtitle Video + Subtitle Video + Subtitle Video-LLaVA [33] 20.3 15.4-4.9 19.8 14.4 -5.4 19.5 16.4-3.1 20.3 16.0-4.3 LLaMA3.2 [21] 27.1 29.3+2.2 27.6 29.6+2.0 25.9 28.1+2.2 27.1 28.8+1.7 Qwen2-VL [64] 31.8 41.1+9.3 30.9 39.4+8.5 34.2 41.8+7.6 32.4 41.2+8.8 mPLUG-Owl3 [75] 33.0 39.2+6.2 32.3 38.3+6.0 34.6 39.2+4.6 32.9 38.7+5.8 LLaVA-OneVision [27] 37.7 44.0+6.3 36.3 42.7+6.4 39.7 45.7+6.0 37.7 43.9+6.2 InternVL2.5 [5] 39.0 48.3+9.3 38.6 47.9+9.3 39.2 47.1+7.9 39.1 47.8+8.7 LLaVA-Video [82] 40.5 45.9+5.4 38.9 44.6+5.7 42.3 47.7+5.4 40.2 45.6+5.4 GPT 4o [25] 42.8 51.1+8.3 40.9 50.2+9.3 43.6 49.9+6.3 42.6 50.1+7.5
@@ Appendix A Quality Control
[…]
@@ Appendix C Evaluation Prompt
Following previous works [17, 31], we adopt the format of “whole video frames + whole subtitles/audios (optional) + question with prompt” as prompt. We show the evaluation prompt across three input configurations: video-only input, video with subtitles, and video with audio content as following.
@@ Appendix D Prompt Sensitivity
Figure 8: Failure Case. We present two error examples.
Appendix E Failure Case
Figure 8 showcases two failure cases of Gemini 1.5 Pro, VideoLLaMA2, and LLaVA-Video. The left case involves a vision understanding and reasoning error, where Gemini 1.5 Pro incorrectly identified the clock time as 10:30 instead of the actual 10:00 displayed, leading to the incorrect answer. This reflects deficiencies in basic visual perception and properly correlating visual information with the question. The right case demonstrates an audio understanding error, where models misjudged the rhythm pattern change in guzheng music, interpreting it as changing from soothing to intense (option D) rather than the correct passionate to soothing (option B). This case indicates that tasks involving interpretation of musical emotion and rhythm patterns remain challenging for existing MLLMs. These cases highlight the current limitations of MLLMs in processing subtle visual details and interpreting subjective audio qualities.
Appendix F Limitation
@@
Appendix G Broader Impacts & Ethics Statement
@@ Appendix H License
Appendix I Datasheets
I.1 Motivation
@@ I.3 Collection Process
I.4 Preprocessing/cleaning/labeling

### v3

@@ 4.3 Roadmap Towards Real-world Understanding
Vision Information. We investigate the impact of visual information through different input configurations: audio-only, audio with video captions, and audio with video frames. As shown in Table 7, visual information generally improves performance, with Gemini 1.5 Pro’s accuracy increasing from $34.6\%$ (audio-only) to $48.0\%$ (+video). However, impact varies across models, with UnifiedIO2 showing inconsistent gains and even degradation with captions.
@@
The results in Table 4 reveal intriguing patterns in how different forms of audio information influence model performance. For Gemini 1.5 Pro, accuracy increases from $34.4\%$ (video-only) to $39.3\%$ with subtitles, and further to $48.0\%$ with original audio. Other models, such as OneLLM and Qwen2.5-Omni, show similar improvements. These results demonstrate that both subtitles and acoustic features (including tone, emotion, and environmental sounds) contribute valuable information, beyond what subtitles alone can capture, emphasizing the importance of complete acoustic cues in omni-modal real-world understanding.
@@
Methods Speech Event Music Overall Video + Subtitle + Audio Video + Subtitle + Audio Video + Subtitle + Audio Video + Subtitle + Audio Unified-IO-2 L (Lu et al., 2024b) 26.8 13.9-12.9 23.1-3.1 26.9 13.5-13.4 23.2-3.7 26.3 15.0-11.3 24.5-1.8 26.6 14.8-11.8 23.3-3.3 Unified-IO-2 XL (Lu et al., 2024b) 25.0 13.0-12.0 24.6-0.4 24.8 12.3-12.5 25.0+0.2 26.7 15.9-10.8 24.7-2.0 25.3 14.1-11.2 24.7-0.6 Unified-IO-2 XXL (Lu et al., 2024b) 27.0 15.6-11.4 25.7-1.3 26.2 14.2-12.0 25.1-1.1 28.4 19.1-9.3 25.9-2.5 27.2 17.2-10.0 25.9-1.3 OneLLM (Han et al., 2024) 12.5 19.6+7.1 23.2+10.7 12.4 19.3+6.9 22.3+9.9 12.4 19.0+6.6 22.5+10.1 12.6 19.6+7.0 22.8+10.2 VideoLLaMA2 Cheng et al. (2024) 17.1 25.5+8.4 25.2+8.1 16.1 24.9+8.8 24.5+8.4 17.7 27.0+9.3 25.6+7.9 17.4 26.1+8.7 25.4+8.0 VITA-1.5 Fu et al. (2025) 37.6 39.1+1.5 36.2-1.4 36.4 38.2+1.8 35.3-1.1 38.7 40.0+1.3 37.1-1.6 37.7 39.3+1.6 36.5-1.2 Qwen2.5-Omni Xu et al. (2025a) 38.7 38.7+0.0 44.8+6.1 37.6 37.7+0.1 43.8+6.2 40.7 40.3-0.4 46.1+5.4 39.2 39.2+0.0 45.2+6.0 Gemini 1.5 Pro (Team et al., 2024a) 34.3 39.6+5.3 49.2+14.9 33.0 38.9+5.9 48.7+15.7 35.4 39.2+3.8 49.8+14.4 34.4 39.3+4.9 48.0+13.6
@@
Methods Speech Event Music Overall Video + Subtitle Video + Subtitle Video + Subtitle Video + Subtitle Video-LLaVA (Lin et al., 2023) 20.3 15.4-4.9 19.8 14.4 -5.4 19.5 16.4-3.1 20.3 16.0-4.3 LLaMA3.2 (Grattafiori et al., 2024) 27.1 29.3+2.2 27.6 29.6+2.0 25.9 28.1+2.2 27.1 28.8+1.7 Qwen2-VL (Wang et al., 2024a) 31.8 41.1+9.3 30.9 39.4+8.5 34.2 41.8+7.6 32.4 41.2+8.8 mPLUG-Owl3 (Ye et al., 2024) 33.0 39.2+6.2 32.3 38.3+6.0 34.6 39.2+4.6 32.9 38.7+5.8 LLaVA-OneVision (Li et al., 2024a) 37.7 44.0+6.3 36.3 42.7+6.4 39.7 45.7+6.0 37.7 43.9+6.2 InternVL2.5 (Chen et al., 2024b) 39.0 48.3+9.3 38.6 47.9+9.3 39.2 47.1+7.9 39.1 47.8+8.7 LLaVA-Video (Zhang et al., 2024d) 40.5 45.9+5.4 38.9 44.6+5.7 42.3 47.7+5.4 40.2 45.6+5.4 GPT 4o (Hurst et al., 2024) 42.8 51.1+8.3 40.9 50.2+9.3 43.6 49.9+6.3 42.6 50.1+7.5
@@
Figure 8: Failure Case. We present two error examples.
Figure 8 showcases two failure cases of Gemini 1.5 Pro, VideoLLaMA2, and LLaVA-Video. The left case involves a vision understanding and reasoning error, where Gemini 1.5 Pro incorrectly identified the clock time as 10:30 instead of the actual 10:00 displayed, leading to the incorrect answer. This reflects deficiencies in basic visual perception and properly correlating visual information with the question. The right case demonstrates an audio understanding error, where models misjudged the rhythm pattern change in guzheng music, interpreting it as changing from soothing to intense (option D) rather than the correct passionate to soothing (option B). This case indicates that tasks involving interpretation of emotion and rhythm patterns remain challenging for existing MLLMs.
@@ Appendix A Quality Control
Sufficient annotation training. We conducted a one-week training program with 200 videos (excluded from final benchmark) until annotators achieved high proficiency (only $10\%$ requiring modifications).
@@ Appendix C Evaluation Prompt
Following previous works (Fu et al., 2024b; Li et al., 2024c), we adopt the format of “whole video frames + whole subtitles/audios (optional) + question with prompt” as prompt. We show the evaluation prompt across three input configurations: video-only input, video with subtitles, and video with audio content as following.
@@ Appendix D Prompt Sensitivity
Appendix E Limitation
@@
Appendix F Broader Impacts & Ethics Statement
@@ Appendix G License
Appendix H Datasheets
H.1 Motivation
@@ H.3 Collection Process
H.4 Preprocessing/cleaning/labeling